# 2TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3 Canada

### v1

@@ Abstract
By extending our analysis and simulations of our 37K $\beta$-decay data set to allow the $\beta$ asymmetry with respect to nuclear spin to vary with $\beta$ energy $E_{\beta}$, we have gained sensitivity to new physics that depends on a helicity factor for the $\beta$, $m_{\beta}/E_{\beta}$. In particular, we constrain Lorentz scalar and tensor quark-lepton interaction strengths at a sensitivity complementary to the similar Fierz interference term in neutron $\beta$ decay. Our result for that new physics is $b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst), consistent with the standard model electroweak interaction value $b_{F}=0$. We consider presently achieved complementarity to $\beta$-decay and particle physics experiments, along with projectable technical improvements to our method.
@@ I Introduction
Here we extend simulations and analysis techniques to allow for different new physics that couples a $\nu$ with weak interaction helicity to a $\beta^{+}$ with the same sign of helicity as the $\nu$, which mutiplies the result by a helicity factor $m_{\beta}/E_{\beta}$ needed to produce that $\beta$ with opposite helicity. Having noted in Ref. [4] that one dominant systematic, $\beta$ scattering off surfaces, strongly depends on $E_{\beta}$ as it approaches $m_{\beta}$, we find that the energy dependence complicates the simulations and analysis considerably.
@@ II Theory and Methods
$A_{\beta}[E]=A_{\beta}/(1+mb/E)\approx A_{\beta}(1-mb/E)$ (2)
We write the approximate expression to emphasize that sensitivity to $b$ in any given decay scales with $A_{\beta}$. The value of $A_{\beta}$ in 37K is reasonably large compared to other well-measured cases like the neutron and 19Ne, making 37K decay more sensitive to $b$ by that simple figure of meri.
We now specialize slightly to the Lee-Yang Lagrangian for electron-nucleon interactions, and call its Fierz interference term explicitly $b_{F}$. It has a Lorentz scalar structure interfering with the standard model vector, and a tensor interfering with axial vector [1]:
$b_{F}=\frac{\pm\gamma[(C_{S}+C_{S}^{\prime})/C_{V}+\rho^{2}(C_{T}+C_{T}^{\prime})/C_{A}]}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+C_{T}^{2}+C_{T}^{\prime 2})/2}$ (3)
@@
$A_{\beta}=\frac{\frac{\pm\rho^{2}}{J+1}(C_{T}C_{T}^{\prime}-C_{A}C_{A}^{\prime})}{1+\rho^{2}+((C_{S}^{2}+C_{S}^{\prime 2})+(C_{T}^{2}+C_{T}^{\prime 2})/2)}$ (4)
$+\frac{\rho\sqrt{\frac{J}{J+1}}(C_{S}C_{T}^{\prime}+C_{S}^{\prime}C_{T}-C_{V}C_{A}^{\prime}-C_{V}^{\prime}C_{A})}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+C_{T}^{2}+C_{T}^{\prime 2})/2}$ (5)
where $\rho$ is the standard generalization for nuclei $\rho=\frac{g_{A}}{g_{V}}\frac{M_{GT}}{M_{F}}$ determined from an average Fermi strength value along with the measured 37K $fT$ value [19]. The $\pm$ signs are for $\beta^{\mp}$ decay.
@@
A 2nd-class term in the axial vector lepton-nucleon current, termed the induced tensor term $\frac{g_{T}}{m}\sigma_{\mu\nu}k_{\nu}$, has different dependence on $E_{\beta}$ than the electron-quark Lorentz tensor interaction [6, 21]. It can be generated by non-standard-model meson exchange and nucleon-nucleon contact interactions [21, 22, 23], and there are tight constraints from tailored nuclear $\beta$ decay observables, in particul the $\beta$-alignment correlation which vanishes in Gamow-Teller decay [12]. In this parameterization of 2nd-class tensor interactions, one term does contribute equally to neutron and nuclear $\beta$ decay, while two other complicated nuclear matrix elements can vary with any nucleus, only one of which vanishes in isobaric mirror decays [22, 23]. Relating constraints on $C_{T}$ is nontrivial, so we don’t comment further below.
@@ III.3 Geometry and Detectors
An electric field collects 37Ar ions produced in 37K $\beta^{+}$ decay, along with photoions from the 355 nm laser in Fig. 1, to an MCP with 78 mm active diameter located 9.7 cm away. By comparing trajectories with a detailed finite element calculation, we found our simulations could assume a uniform 150 V/cm field. The photoion information hleps determine the cloud polarization, center in space, and spatial distribution. The 37Ar ion info is not used in this paper’s analysis.
Decay by $\beta^{+}$ naturally feeds the negative Ar-1 ion. We constrained contributions from the known Ar-1 $\tau$=250 ns metastable by the lack of a significant tail on the timing spectrum between shakeoff electron and $\beta$: nevertheless, to eliminate possible bias on $A_{\beta}$ from loss of detection efficiency from the delayed shakeoff emission and its dependence on the recoil energy and direction, we don’t impose a vertical position cut on the shakeoff electron detector [4].
@@ III.4 Data and interpretation
We found it important in Ref. [4] to remove backgrounds in our geometry from decays of untrapped atoms by requiring the $\beta^{+}$ to be in coincidence with shakeoff atomic electrons (SOE’s). To improve our knowledge of the asymmetry at low $E_{\beta}$ to study $b_{F}$, we found the relative timing to be even more important. So we removed a residual timing walk in the scintillator, produced at low $E_{\beta}$ pulse height, which had not been removed by the constant fraction discriminator. The resulting improved timing (Fig. 2) nearly removes a small background from untrapped atoms on surfaces that is more apparent at low $E_{\beta}$. The simulated background is scaled to a tail of the timing spectrum, as we do not know what fraction of untrapped atoms sticks long enough to decay. This background was one of our larger systematics in Ref. [4], while with the improved timing the correction from the simulation shown is nearly negligible and introduces small uncertainty. Also, we can now understand and model well the measured timing signal from the known cloud size using a combination of 9% very low-energy electrons– probably Ar- autoionization– 77% 4S electron shakeoff characterized by their 4.5 eV binding energy, and 14% 3P electron shakeoff characterized by their 18 eV binding energy. The SOE energy spectra use the sudden approximation calculations with hydrogenic wavefunctions by Levinger [11].
@@ III.5 Extraction of $A_{\beta}[E_{\beta}]$
We extract $A_{\beta}$ and $b_{F}$ from this info. We calculate $\chi^{2}$ from the difference between the superratio of all data and simulations with varied $A_{\beta}$ and $b_{F}$.
[…]
To extract our two degrees of freedom, we then look for this total $\chi^{2}$ to increase from its minimum by 2.3 when varying $A_{\beta}$ and $b_{F}$ [13]. The result, shown in Fig. 4, is $b_{F}=-0.0002\pm 0.024$ (statistical). We consider systematics in the next section.
Figure 4: A $\chi^{2}$ map of all experimental data compared to a simulated parameter space of $A_{\beta}$ and $b$, showing in solid lines the 1 $\sigma$ 68% confidence limits from a change in $\chi^{2}$ of 2.3 from the minimum. Dashed lines are 90% CL. The tighter, black ellipses add the $chi^{2}$ contribution from the predicted center value of $A_{\beta}$, which determines $b_{F}$ of Eq. 3 from electron-quark Lorentz scalar and tensor intereactions. The larger, red ellipses determine the more general measurement of $b$ from Eq. 1 without the FtFt constraint.
III.5.1 Peripheral quantities
For completeness, we consider two other extremes to extract information from our data set.
As a check, by forcing $b_{F}$=0, i.e. introducing only the same energy-independent physics allowed by our Ref. [4], we deduce $A_{\beta}$= -0.5712 $\pm$ 0.013 (stat). Since forcing $b_{F}$=0 produces a 1-parameter fit, we read off the change in $\chi^{2}$ by 1, not the 2.3 change for the 2-parameter fits. This would imply that our reanalysis in this paper, primarily the time walk experimental correction of Fig. 2, would alter our previous answer for the $\beta$ asymmetry with no extra energy dependence from $A_{\beta}=-0.5707(13)(13)$ to –0.5713(13)(10). The smaller systematic uncertainty is from the near-elimination of the background from untrapped atoms. The centroid change is about one-third of our total uncertainty, so the upper limits on right-handed currents in Ref. [4] and the extraction of $V_{ud}$ are essentially unchanged, and we do not comment further.
Although the Lee-Yang Lagrangian is the only model we are aware of for $b$, we also consider a more general situation where $b$ and $A_{\beta}$ are allowed to vary independently without the prediction of $A_{\beta}$ from the 37K $fT$ value. E.g, this would happen if V+A, S, and T currents were all allowed to vary independently. It should be intuitive that there are large correlations between $A_{\beta}$ and $b$ in Fig. 3, so the result inflates the uncertainty on both: $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.0039 (syst) and $A_{\beta}=-0.5738\pm 0.0082{\rm(stat)}\pm 0.0041{\rm(syst)}$. We re-state for emphasis that these are not primary results of this paper.
@@ III.6 Systematic uncertainties
Table 1: Systematic uncertainty budget for the two-parameter analysis for $b$ (which also applies to $b_{F}$) and $A_{\beta}$. All uncertainties are believed to be uncorrelated, and are added in quadrature. See text for future projections
@@ IV Results
Our result in the Lee-Yang scalar-tensor Lagrangian of quark-lepton non-SM Lorentz current is then
$b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst),
We explore complementarity of this our result to other experiments next.
@@ IV.1 Sensitivity to new physics
We show a 90% CL (i.e. 1.65 $\sigma$) exclusion plot of Lorentz scalar vs. tensor lepton-quark couplings in Fig. 5, assuming no coupling to non-standard model right-handed neutrinos. There exist complete reviews setting constraints on the Lee-Yang Lagrangian parameters [14], and separately reviews on 2nd-class currents [21, 12]. To allow for the possibility of non-universal interactions, here we show ten $\beta$ decay and other particle physics constraints, without review or averaging. Our goal is to show complementarity in situations with less universal physics models than the Lee-Yang Lagrangian.
Figure 5: Exclusion plot of Lorentz scalar and tensor lepton-quark couplings, 90% CL. Consistency between neutron $\beta$ decay, nuclear $\beta$ decay, and particle physics requires a non-universal interaction with scale less than TeV. See details and references in text.
Our 37K results could be said to complement the neutron results on $C_{S}$ because our smaller Gamow-Teller strength weighs $C_{S}$ more than $C_{T}$, yet our own group’s direct constraints on $C_{S}$ from 38mK decay [26] are considerably tighter, shown here averaged with the 32Ar experiment with similar sensitivity [24, 25].
@@
The other tight indirect constraints on $C_{S}$, shown in the green vertical band, are from the consistency of $0^{+}\rightarrow 0^{+}$ $Ft[E_{\beta}]$ values in a 2020 global survey [31]. Yet allowing $b_{F}$ to float instead of fixing it to zero would inflate the uncertainty in $V_{ud}$ by factors of two, motivating independent determinations of $b_{F}$ with similar sensitivity. Recent theory developments may relax or improve the sensitivity: re-evaluation of the lepton phase space $f$ summarized in Ref. [50] with better calculation of a radiative correction in 10C decay [51]; continuing theoretical challenges to support existing isospin-breaking calculations, including a possible need for first-principles inclusion of strong interaction isospin breaking [52] rather than fitting the Nolen-Schiffer anomaly [31]. Note that a fit to Ref. [31] of the 2nd-class induced scalar contribution to an electron-nucleon vector current parameterized as $\frac{e}{A}\frac{m_{\beta}}{m_{\rm nucleon}}\frac{m_{\beta}}{E_{\beta}}$ [54, 20], produces $e=-4\pm 32$.
Two experiments are shown for neutron $\beta$ decay. Measurement of $\beta$ decay asymmetry and spectrum shape has competitive sensitivity to $b_{F}$ consistent with zero (solid black lines) [32]. An analysis including the recoil energy spectrum of protons and Ref. [32] data together, but floating $g_{A}/g_{V}$ to a different result from either, asserts a 2.8 $\sigma$ nonzero $b_{F}$ result (the dashed black line) [33]. By inspection, the positive result of Ref. [33] can be reconciled with the tight constraint on Lorentz tensor lepton-quark interactions from $\pi\rightarrow e\nu\gamma$ decay [39] with a nonzero and negative $C_{S}$ of -0.05. There are by inspection varying degrees of tension with the above-mentioned Lorentz electron-quark scalar current constraints [26, 24, 25, 31], with isobaric mirror decays like the present 37K work, with 19Ne (from Ref. [35] which used its $fT$ measurement together with $A_{\beta}[E_{\beta}]$ [34], both under revision [36]), and with the relative helicity of $\beta$’s in Fermi vs. Gamow-Teller decay (the cyan band denoted $P_{F}/P_{GT}$[38]. Neverthless, we note that interpreting the $C_{S}\approx$ -0.05 region as an induced scalar for neutron decay $e\approx$ 31 alleviates tension with all other nuclear $\beta$ decay because of its $1/A$ dependence.
We see from Fig. 5 that to be complementary to other physics, we would have to improve our 37K $b_{F}$ result to compete with the best isospin mirror nucleus results in 19Ne [35, 34, 36]; measurements of $a_{\beta\nu}$ in 21Na [37] would need similar improvement. We hope demonstrated and in-progress improvements in the Experiment suggestions will let us achieve the factor of five needed.

### v4

@@ Abstract
By extending our analysis and simulations of our 37K $\beta$-decay data set to allow the $\beta$ asymmetry with respect to nuclear spin to vary with $\beta$ energy $E_{\beta}$, we have gained sensitivity to new physics that depends on a helicity factor for the $\beta$, $m_{\beta}/E_{\beta}$. In particular, we constrain Lorentz scalar and tensor quark-lepton interaction strengths at a sensitivity complementary to the similar Fierz interference term in neutron $\beta$ decay. Our result for that new physics is a Fierz interference term $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.039 (syst), consistent with the standard model electroweak interaction value $b=0$. We consider presently achieved complementarity to $\beta$-decay and particle physics experiments, along with projectable technical improvements to our method.
@@ I Introduction
Here we extend simulations and analysis techniques to allow for different new physics that couples a $\nu$ with weak interaction helicity to a $\beta^{+}$ with the same sign of helicity as the $\nu$, which multiplies the result by a helicity factor $m_{\beta}/E_{\beta}$ needed to produce that $\beta$ with opposite helicity. Having noted in Ref. [4] that one dominant systematic, $\beta$ scattering off surfaces, strongly depends on $E_{\beta}$ as it approaches $m_{\beta}$, we find that the energy dependence complicates the simulations and analysis considerably.
@@ II Theory and Methods
$A_{\beta}[E]=A_{\beta}/(1+bm/E)\approx A_{\beta}(1-bm/E)$ (2)
We write the approximate expression to emphasize that sensitivity to $b$ in any given decay scales with $A_{\beta}$. The value of $A_{\beta}$ in 37K is reasonably large compared to other well-measured cases like the neutron and 19Ne, making 37K decay more sensitive to $b$ by that simple figure of merit.
In the Lee-Yang Lagrangian, $b$ is produced by a Lorentz scalar interaction interfering with the standard model Lorentz vector interaction, and a Lorentz tensor interaction interfering with a Lorentz axial vector interaction [1]:
$b=\frac{\pm\gamma[(C_{S}+C_{S}^{\prime})+\rho^{2}(C_{T}+C_{T}^{\prime})]}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+\rho^{2}(C_{T}^{2}+C_{T}^{\prime 2}))/2}$ (3)
@@
$A_{\beta}=\frac{\frac{\pm\rho^{2}}{J+1}(C_{T}C_{T}^{\prime}-1)}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+\rho^{2}(C_{T}^{2}+C_{T}^{\prime 2}))/2}$
$+\frac{\rho\sqrt{\frac{J}{J+1}}(C_{S}C_{T}^{\prime}+C_{S}^{\prime}C_{T}-2)}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+\rho^{2}(C_{T}^{2}+C_{T}^{\prime 2}))/2}$ (4)
where $\rho=\frac{g_{A}}{g_{V}}\frac{M_{GT}}{M_{F}}$ is determined by us below from our collaboration’s measured 37K $fT$ value [19], together with an average $M_{F}$ from $0^{+}\rightarrow 0^{+}$ decays [32]. The $\pm$ signs are for $\beta^{\mp}$ decay.
@@
A 2nd-class term in the axial vector lepton-nucleon current, termed the induced tensor term $\frac{g_{T}}{m_{\rm nucleon}}\sigma_{\mu\nu}k_{\nu}$, has different dependence on $E_{\beta}$ than the electron-quark Lorentz tensor interaction [6, 21]. It can be generated by non-standard-model meson exchange and nucleon-nucleon contact interactions [21, 22, 23], and there are tight constraints from tailored nuclear $\beta$ decay observables, in particular the $\beta$-alignment correlation which vanishes in Gamow-Teller decay [12]. In this parameterization of 2nd-class tensor interactions, one term does contribute equally to neutron and nuclear $\beta$ decay, while two other complicated nuclear matrix elements can vary with any nucleus, only one of which vanishes in isobaric mirror decays [22, 23]. Relating constraints to those on $C_{T}$ is nontrivial, so we don’t comment further below.
@@ III.3 Geometry and Detectors
An electric field collects 37Ar ions produced in 37K $\beta^{+}$ decay, along with photoions from the 355 nm laser in Fig. 1, to an MCP with 78 mm active diameter located 9.7 cm away. By comparing trajectories with a detailed finite element calculation, we found our simulations could assume a uniform 150 V/cm field. The photoion information helps determine the cloud polarization, center in space, and spatial distribution. The 37Ar ion info is not used in this paper’s analysis.
Decay by $\beta^{+}$ naturally feeds the negative Ar-1 ion. We constrained contributions from the known Ar-1 $\tau$=250 ns metastable state by the lack of a significant tail on the timing spectrum between shakeoff electron and $\beta$: nevertheless, to eliminate possible bias on $A_{\beta}$ from loss of detection efficiency from the delayed shakeoff emission and its dependence on the recoil energy and direction, we don’t impose a vertical position cut on the shakeoff electron detector [4].
@@ III.4 Data and interpretation
We found it important in Ref. [4] to remove backgrounds in our geometry from decays of untrapped atoms by requiring the $\beta^{+}$ to be in coincidence with shakeoff atomic electrons (SOE’s). To improve our knowledge of the asymmetry at low $E_{\beta}$ to study $b$, we found the relative timing to be even more important. So we removed a residual timing walk in the scintillator, produced at low $E_{\beta}$ pulse height, which had not been removed by the constant fraction discriminator. The resulting improved timing (Fig. 2) nearly removes a small background from untrapped atoms on surfaces that is more apparent at low $E_{\beta}$. The simulated background is scaled to a tail of the timing spectrum, as we do not know what fraction of untrapped atoms sticks long enough to decay. This background was one of our larger systematics in Ref. [4], while with the improved timing the correction from the simulation shown is nearly negligible and introduces small uncertainty. Also, we can now understand and model well the measured timing signal from the known cloud size using a combination of 9% very low-energy electrons– probably Ar- autoionization– 77% 4S electron shakeoff characterized by their 4.5 eV binding energy, and 14% 3P electron shakeoff characterized by their 18 eV binding energy. The SOE energy spectra use the sudden approximation calculations with hydrogenic wavefunctions by Levinger [11].
@@ III.5 Extraction of $A_{\beta}[E_{\beta}]$
We extract $A_{\beta}$ and $b$ from this info. We calculate $\chi^{2}$ from the difference between the superratio of all data and simulations with varied $A_{\beta}$ and $b$. To extract our two degrees of freedom, we then look for this total $\chi^{2}$ to increase from its minimum by 2.3 when varying $A_{\beta}$ and $b$ [13]. Allowing $b$ and $A_{\beta}$ to vary independently produces our main result, $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.0039 (syst) and $A_{\beta}=-0.5738\pm 0.0082{\rm(stat)}\pm 0.0041{\rm(syst)}$. The linear correlation between $A_{\beta}$ and $b$ reflected in Fig. 4 can be expressed by $A_{\beta}=-0.5707-0.090*b$.
We find more sensitivity to new physics by adding information. We also show the standard model prediction for $A_{\beta}=-0.5706$ in Fig. 4, given $\rho$ and its uncertainty from the measured 37K $Ft$ of our collaboration [19], and taking $M_{F}$ from the measured $0^{+}\rightarrow 0^{+}$ average $Ft_{0}$ value [32]. In Fig. 4 we are not considering the variation in $Ft_{0}$ with $C_{S}+C^{\prime}_{S}$, as we wish to show 37K complementarity to such observables below. We further show $\chi^{2}$ contours as black ellipses in Fig. 4 including the contribution from the experimental uncertainty on 37K $Ft$. When Lorentz scalar and tensor interactions are considered below, we correlate $b$ variations with changes in $A_{\beta}$ by Eqs. 3-4, and plot the resulting combined constraint on those couplings in Fig. 5.
Figure 4: A $\chi^{2}$ map of all experimental data compared to a simulated parameter space of $A_{\beta}$ and $b$, showing in solid lines the 1 $\sigma$ 68% confidence limits from a change in $\chi^{2}$ of 2.3 from the minimum. Dashed lines are 90% CL. The larger, red ellipses determine the more general measurement of $b$ from Eq. 1. The tighter, black ellipses add the $chi^{2}$ contribution from the predicted center value of $A_{\beta}$ from FtFt information (see text), which determines $C_{S}+C^{\prime}_{S}$ vs. $C_{T}+C^{\prime}_{T}$ in Fig. 5 below.
III.5.1 Check of previous result
As a check, by forcing bb=0, i.e. introducing only the same energy-independent physics allowed by our Ref. [4], we deduce $A_{\beta}$= -0.5712 $\pm$ 0.013 (stat). Since forcing bb=0 produces a 1-parameter fit, we read off the change in $\chi^{2}$ by 1, not the 2.3 change for the 2-parameter fits. This deductions implies that our reanalysis in this paper, primarily the time walk experimental correction of Fig. 2, alters our previous answer for the $\beta$ asymmetry with no extra energy dependence from $A_{\beta}=-0.5707(13)(13)$ to -0.5712(13)(10). The smaller systematic uncertainty is from the near-elimination of the background from untrapped atoms. The centroid change is about one-third of our total uncertainty, so the upper limits on right-handed currents in Ref. [4] and the extraction of $V_{ud}$ are essentially unchanged, and we do not comment further.
@@ III.6 Systematic uncertainties
Table 1: Systematic uncertainty budget from the two-parameter analysis for $b$.All uncertainties are believed to be uncorrelated, and are added in quadrature. For details on future projections, see text.
@@ IV Results
Our result for the Fierz interference term is then $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.039 (syst). We plot this 1 $\sigma$ constraint, using Eq. 3 only, on Fig. 5 (the thick black lines).
We also plot (the large black ellipse) our 1 $\sigma$ allowed area considering, as mentioned above, our present work’s measured $A_{\beta}$ and $b$, their correlation through Eqs. 3–4, our collaboration’s measured 37K $Ft$ value [19], and the average $Ft$ value of $0^{+}\rightarrow 0^{+}$ decays [32]. This produces a finite constraint on Lorentz tensor couplings and a considerably tighter constraint on Lorentz scalar couplings in Fig. 5.
We explore complementarity of our results to other experiments next.
@@ IV.1 Sensitivity to new physics
We show a 1 $\sigma$ exclusion plot of Lorentz scalar vs. tensor lepton-quark couplings in Fig. 5, assuming no coupling to non-standard model right-handed neutrinos. There exist complete reviews setting constraints on the Lee-Yang Lagrangian parameters [14], and separately reviews on 2nd-class currents [21, 12]. To allow for the possibility of non-universal interactions, here we show ten $\beta$ decay and other particle physics constraints, without review or averaging. Our goal is to show complementarity in situations with less universal physics models than the Lee-Yang Lagrangian.
Figure 5: Exclusion plot at 1 $\sigma$ of Lorentz scalar and tensor lepton-quark couplings to left-handed $\nu$’s and wrong-handed $\beta$’s. Consistency between neutron $\beta$ decay, nuclear $\beta$ decay, and particle physics requires a non-universal interaction with scale less than TeV. Our present 37K constraints from $b$ (solid black lines), along with our tighter constraints combining $b$, $A_{\beta}$ and fTfT info (black ellipse), are also detailed in text.
Our 37K results could be said to complement the neutron results on $C_{S}$, because our smaller Gamow-Teller strength weighs $C_{S}$ more than $C_{T}$. Yet our own group’s direct constraints on $C_{S}$ from 38mK decay [27] are tighter, shown here averaged with the 32Ar experiment with similar sensitivity [25, 26].
@@
The other tight indirect constraints on $C_{S}$, shown in the green vertical band, are from the consistency of $0^{+}\rightarrow 0^{+}$ $Ft[E_{\beta}]$ values in a 2020 global survey [32]. Yet allowing $b$ to float instead of fixing it to zero is known to inflate the uncertainty in $V_{ud}$ by about two [14], motivating independent determinations of $b$ with similar sensitivity. Recent theory developments may relax or improve the sensitivity: re-evaluation of the lepton phase space $f$ summarized in Ref. [51] with better calculation of a radiative correction in 10C decay [52]; continuing theoretical challenges to support existing isospin-breaking calculations, including a possible need for first-principles inclusion of strong interaction isospin breaking [53] rather than fitting the Nolen-Schiffer anomaly [32]. Note that a fit to Ref. [32] of the 2nd-class induced scalar contribution to an electron-nucleon vector current parameterized as $\frac{e}{A}\frac{m_{\beta}}{m_{\rm nucleon}}\frac{m_{\beta}}{E_{\beta}}$ [55, 20], produces $e=-4\pm 32$.
Two experiments are shown for neutron $\beta$ decay. Measurement of $\beta$ decay asymmetry and spectrum shape has competitive sensitivity to $b$ consistent with zero (solid red lines) [33]. An analysis including the recoil energy spectrum of protons and Ref. [33] data together, but floating $g_{A}/g_{V}$ to a different result from either, asserts a 2.8 $\sigma$ nonzero $b$ result (the dashed red line) [34]. By inspection, the positive result of Ref. [34] can be reconciled with the tight constraint on Lorentz tensor lepton-quark interactions from $\pi\rightarrow e\nu\gamma$ decay [40] with a nonzero and negative $C_{S}$ of -0.05. There are by inspection varying degrees of tension with the above-mentioned Lorentz electron-quark scalar current constraints [27, 25, 26, 32], with isobaric mirror decays like the present 37K work, with 19Ne (from Ref. [36] which used its $fT$ measurement together with $A_{\beta}[E_{\beta}]$ [35], both under revision [37]), and with the relative helicity of $\beta$’s in Fermi vs. Gamow-Teller decay (the cyan band denoted $P_{F}/P_{GT}$[39]. Neverthless, we note that interpreting the $C_{S}\approx$ -0.05 region as an hypothetical induced scalar for neutron decay with constant form factor $e\approx$ 31 alleviates tension with all other nuclear $\beta$ decay because of the resulting $1/A$ dependence of its contribution.
We see from Fig. 5 that to be complementary to other physics, we would have to improve our 37K $b$ result to compete with the best isospin mirror nucleus results in 19Ne [36, 35, 37] and to other measurements. As we are presently dominated by statistics, and project to improve systematics, we hope our demonstrated and in-progress improvements in the experiment will let us achieve the order of magnitude needed.